# The Computational Complexity of Counting Linear Regions in ReLU Neural Networks

### v2

@@ B.1 Omitted proofs for the one hidden layer case
Let NN be a ReLU network with one hidden layer and let $n_{1}]$ denote the set of neurons. For ease of notation, we denote $A^{(2)}=(a_{1},\dots,a_{n_{1}})\in\mathbb{R}^{1\times n_{1}}$ and $A^{(1)}=(w_{ij})_{i\in[n_{1}],j\in[n]}$. Then, the function computed by the network is
@@
We only consider Definitions˜3, 4, 5 and 6 here (for Definitions˜1 and 2, the answer is trivially yes, as $n_{1}>0$). Let NN be a ReLU network as defined above.
The idea is to check for every neuron if the function $f_{N}$ computed by the ReLU network NN has breakpoints along the hyperplane corresponding to the neuron, that is, we check if the gradient of $f_{N}$ is discontinuous along the hyperplane. Since multiple neurons can correspond to the same hyperplane, it is possible that the functions of these neurons cancel such that no breakpoints along the hyperplane are introduced. First, we group the neurons according to the hyperplanes to which they correspond.
@@
the sum of all functions having $H_{1}$ as corresponding hyperplane is affine and $f_{N}$ has no breakpoints along the hyperplane $H_{1}$. Otherwise, $f_{N}$ has breakpoints along the hyperplane $H_{1}$ and NN has more than one linear region.
Thus, NN has only a single linear region if and only if (1) holds for all $j\in[n_{1}]$ (replace 1 by $j$ in (1) and define the set $I_{j}$ and constants $c_{ji}$ as before with $j$ instead of 1), which can be verified in polynomial time. ∎
@@
A certificate for Definitions˜1 and 2 is simply an activation pattern $a\in\{0,1\}^{s(N)}$ of the ReLU network NN, which can be checked in polynomial time by computing the dimension of $S_{a}$, see Lemma˜A.5. Thus, Linear region counting is in $\mathsf{P}$ for Definitions˜1 and 2.
We now construct certificates for Definitions˜3 and 4. Given a ReLU network NN with one hidden layer, the set $\mathcal{H}=(H_{i})_{i\in[m]}$ of hyperplanes that correspond to breakpoints of the function $f_{N}$ can be computed in polynomial time using the procedure described in the proof of Theorem˜4.1. By construction, the function $f_{N}$ is affine on each cell of the hyperplane arrangement $\mathcal{H}$ and the affine functions that are realized on two neighboring cells (cells with an (n1)(n-1)-dimensional intersection) cannot be equal. Thus, each cell of the hyperplane arrangement $\mathcal{H}$ is an open connected region.
Since each open connected region is convex, the set of convex regions of the ReLU network NN is well defined and its cardinality is equal to the set of open connected regions.
@@ B.2 Omitted proofs for more than one hidden layers
Intuition for the proof of Lemma˜5.3. The proof of Lemma˜5.3 improves the reduction of Wang [2022], who use a result of Katz et al. [2017, Appendix I]. Similar to Katz et al. [2017, Appendix I], we rely on the simple fact that given a $\mathsf{SAT}$ formula $\phi(x)=\bigwedge_{i=1}^{m}((\bigvee_{j\in J^{+}_{i}}x_{j})\vee(\bigvee_{j\in J^{-}_{i}}\neg x_{j})$ on the variables $x_{1},\dots,x_{n}$, the function
@@
We first show that the problem is in $\mathsf{NP}$. If a ReLU network NN is a yes-instance of 1-region-decision, then there are two proper activation regions on which two distinct affine functions are computed. Two activation patterns corresponding to such proper activation regions serve as a polynomial certificate of a yes-instance. Given two vectors $a,a^{\prime}\in\{0,1\}^{s(N)}$, we can verify the certificate in polynomial time. First, we check if $a$ and $a^{\prime}$ correspond to proper activation regions by computing the dimension of $S_{a}$ and $S_{a^{\prime}}$ in polynomial time, see Lemma˜A.5. If $S_{a}$ and $S_{a^{\prime}}$ are proper activation regions and $f_{N}^{a}\neq f_{N}^{a^{\prime}}$ holds, which can be checked in polynomial time (see Lemma˜A.4), then $a,a^{\prime}$ is a valid certificate of a yes-instance. Thus, the problem is in $\mathsf{NP}$.
@@
Let $L\in\mathbb{N}$ be a fixed constant. Given two ReLU networks $N,N^{\prime}$ with $L$ hidden layers, let $N^{-}$ represent the network with $L$ hidden layers that ‘subtracts’ $N^{\prime}$ from NN by computing the networks NN and $N^{\prime}$ in parallel. NN and $N^{\prime}$ compute the same function if and only if $N^{-}$ computes the zero function.
@@
Given a subset II of $[n]$, we define the set
@@
Let $N_{\phi}^{(k)}$ be the ReLU network composed of taking $k$ copies of NN each with a disjoint set of $l$ variables. The function computed by the ReLU network $N^{(k)}$ is $f_{N_{\phi}^{(k)}}:\mathbb{R}^{lk}\to\mathbb{R}$ with
@@ B.3 Omitted proofs for polynomial space algorithms
Given a ReLU network NN and an affine map $\varphi(x_{1},...,x_{n})=\sum_{i=1}^{n}a_{i}x_{i}+b$, we can check in space polynomial in $\langle N\rangle$ and in the encoding size of the coefficients of $\varphi$ whether $\varphi$ is the function of an affine region of NN.
@@
First, note that $\varphi$ is the function of an affine region of NN if and only if there is a proper activation region on which $\varphi$ is realized.
Now, go over all $2^{s(N)}$ possible activation patterns for neurons of NN using space polynomial in $s(N)$. For each vector $a\in\{0,1\}^{s(n)}$, it is possible to verify in time polynomial in the encoding size $\langle N\rangle$ of the ReLU network NN whether $S_{a}$ is a proper activation region, see Lemma˜A.5. Further, we can check in polynomial time whether $\varphi$ is equal to the function $f^{a}_{N}$ computed on the proper activation region $S_{a}$, see Lemma˜A.4. ∎
@@
Let NN be a ReLU network. As discussed in Section˜6, the number of activation regions and proper activation regions can be counted in space which is polynomial in the encoding size $\langle N\rangle$ of the ReLU network NN. Now, we describe a polynomial space algorithm for counting the number of affine regions.
By Lemma˜A.3, the encoding size of any coefficient of an affine function that occurs in one of the affine regions of NN is bounded by $M:=36d^{2}n_{\max}^{2}\langle A_{\max}\rangle$, which is polynomial in $\langle N\rangle$.
As each coefficient of an affine function of NN is a fraction, $M$ is also an upper bound on the encoding size of a numerator and on the encoding size of a denominator. Since each affine function of NN is defined by $n+1$ fractions, we can exhaustively search through all sequences of $n+1$ fractions, where the numerator and denominator of each fraction can have encoding size of at most $M$. For each sequence of $n+1$ fractions, by Lemma˜B.4 we can compute in space that is polynomial in $\langle N\rangle$ if the corresponding affine function is the function of an affine region of NN. If an affine function of an affine region is found, we increase a counter by 1. To avoid counting the same affine function more than once, we only check fraction sequences win which the numerator and denominator of every fraction are relatively prime. ∎
@@ C.1 A closed connected region which is not a closure of a union of a set of activation regions
We turn this function into a ReLU network NN with three hidden layers, as illustrated in Figure˜5. For the orange closed connected region $P$ in Figure˜5, there is no set of activation regions such that $P$ is the closure of a union of activation regions of the ReLU network NN.
Figure 5: A ReLU network NN computing $\min(y,\max(-1,-x),\max(3-2x,-x))$. An activation region of NN is either a blue line, blue point, or a full dimensional cell as defined by the blue lines. There are four closed connected region as indicated by the colors. The line between the points (1,1)(1,-1) and (3,3)(3,-3) belongs to the green as well as the orange region.
@@ C.2 Further examples
Figure 6: The functions $g_{\varphi}(x)=1-\max(0,x_{1})-\max(0,1-x_{1}-x_{2})$ (top), $h_{\varphi,\varepsilon}(x)=\max(0,\varepsilon-1+g_{\varepsilon}(x)$ (center), and $f_{N_{\varphi}}(x)=\max(0,T_{2}(x)+\varepsilon-1+g_{\varphi}(x))$ (bottom) for $\varepsilon=0.4$. The function $f_{N_{\varphi}}$ is only nonzero in the blue region, which is contained in the $\varepsilon$-square (orange) around the only satisfying point of $\varphi$ (black).

### v3

@@ B.1 Omitted proofs for the one hidden layer case
Let $N$ be a ReLU network with one hidden layer and let $n_{1}]$ denote the set of neurons. For ease of notation, we denote $A^{(2)}=(a_{1},\dots,a_{n_{1}})\in\mathbb{R}^{1\times n_{1}}$ and $A^{(1)}=(w_{ij})_{i\in[n_{1}],j\in[n]}$. Then, the function computed by the network is
@@
We only consider Definitions˜3, 4, 5 and 6 here (for Definitions˜1 and 2, the answer is trivially yes, as $n_{1}>0$). Let $N$ be a ReLU network as defined above.
The idea is to check for every neuron if the function $f_{N}$ computed by the ReLU network $N$ has breakpoints along the hyperplane corresponding to the neuron, that is, we check if the gradient of $f_{N}$ is discontinuous along the hyperplane. Since multiple neurons can correspond to the same hyperplane, it is possible that the functions of these neurons cancel such that no breakpoints along the hyperplane are introduced. First, we group the neurons according to the hyperplanes to which they correspond.
@@
the sum of all functions having $H_{1}$ as corresponding hyperplane is affine and $f_{N}$ has no breakpoints along the hyperplane $H_{1}$. Otherwise, $f_{N}$ has breakpoints along the hyperplane $H_{1}$ and $N$ has more than one linear region.
Thus, $N$ has only a single linear region if and only if (1) holds for all $j\in[n_{1}]$ (replace 1 by $j$ in (1) and define the set $I_{j}$ and constants $c_{ji}$ as before with $j$ instead of 1), which can be verified in polynomial time. ∎
@@
A certificate for Definitions˜1 and 2 is simply an activation pattern $a\in\{0,1\}^{s(N)}$ of the ReLU network $N$, which can be checked in polynomial time by computing the dimension of $S_{a}$, see Lemma˜A.5. Thus, Linear region counting is in $\mathsf{P}$ for Definitions˜1 and 2.
We now construct certificates for Definitions˜3 and 4. Given a ReLU network $N$ with one hidden layer, the set $\mathcal{H}=(H_{i})_{i\in[m]}$ of hyperplanes that correspond to breakpoints of the function $f_{N}$ can be computed in polynomial time using the procedure described in the proof of Theorem˜4.1. By construction, the function $f_{N}$ is affine on each cell of the hyperplane arrangement $\mathcal{H}$ and the affine functions that are realized on two neighboring cells (cells with an (n1)(n-1)-dimensional intersection) cannot be equal. Thus, each cell of the hyperplane arrangement $\mathcal{H}$ is an open connected region.
Since each open connected region is convex, the set of convex regions of the ReLU network $N$ is well defined and its cardinality is equal to the set of open connected regions.
@@ B.2 Omitted proofs for more than one hidden layers
Intuition for the proof of Lemma˜5.3. The proof of Lemma˜5.3 improves the reduction of Wang [2022], who use a result of Katz et al. [2017, Appendix I]. Similar to Katz et al. [2017, Appendix I], we rely on the simple fact that given a $\mathsf{SAT}$ formula $\phi(x)=\bigwedge_{i=1}^{m}((\bigvee_{j\in J^{+}_{i}}x_{j})\vee(\bigvee_{j\in J^{-}_{i}}\neg x_{j}))$ on the variables $x_{1},\dots,x_{n}$, the function
@@
We first show that the problem is in $\mathsf{NP}$. If a ReLU network $N$ is a yes-instance of 1-region-decision, then there are two proper activation regions on which two distinct affine functions are computed. Two activation patterns corresponding to such proper activation regions serve as a polynomial certificate of a yes-instance. Given two vectors $a,a^{\prime}\in\{0,1\}^{s(N)}$, we can verify the certificate in polynomial time. First, we check if $a$ and $a^{\prime}$ correspond to proper activation regions by computing the dimension of $S_{a}$ and $S_{a^{\prime}}$ in polynomial time, see Lemma˜A.5. If $S_{a}$ and $S_{a^{\prime}}$ are proper activation regions and $f_{N}^{a}\neq f_{N}^{a^{\prime}}$ holds, which can be checked in polynomial time (see Lemma˜A.4), then $a,a^{\prime}$ is a valid certificate of a yes-instance. Thus, the problem is in $\mathsf{NP}$.
@@
Let $L\in\mathbb{N}$ be a fixed constant. Given two ReLU networks $N,N^{\prime}$ with $L$ hidden layers, let $N^{-}$ represent the network with $L$ hidden layers that ‘subtracts’ $N^{\prime}$ from $N$ by computing the networks $N$ and $N^{\prime}$ in parallel. $N$ and $N^{\prime}$ compute the same function if and only if $N^{-}$ computes the zero function.
@@
Given a subset $I$ of $[n]$, we define the set
@@
Let $N_{\phi}^{(k)}$ be the ReLU network composed of taking $k$ copies of $N$ each with a disjoint set of $l$ variables. The function computed by the ReLU network $N^{(k)}$ is $f_{N_{\phi}^{(k)}}:\mathbb{R}^{lk}\to\mathbb{R}$ with
@@ B.3 Omitted proofs for polynomial space algorithms
Given a ReLU network $N$ and an affine map $\varphi(x_{1},...,x_{n})=\sum_{i=1}^{n}a_{i}x_{i}+b$, we can check in space polynomial in $\langle N\rangle$ and in the encoding size of the coefficients of $\varphi$ whether $\varphi$ is the function of an affine region of $N$.
@@
First, note that $\varphi$ is the function of an affine region of $N$ if and only if there is a proper activation region on which $\varphi$ is realized.
Now, go over all $2^{s(N)}$ possible activation patterns for neurons of $N$ using space polynomial in $s(N)$. For each vector $a\in\{0,1\}^{s(n)}$, it is possible to verify in time polynomial in the encoding size $\langle N\rangle$ of the ReLU network $N$ whether $S_{a}$ is a proper activation region, see Lemma˜A.5. Further, we can check in polynomial time whether $\varphi$ is equal to the function $f^{a}_{N}$ computed on the proper activation region $S_{a}$, see Lemma˜A.4. ∎
@@
Let $N$ be a ReLU network. As discussed in Section˜6, the number of activation regions and proper activation regions can be counted in space which is polynomial in the encoding size $\langle N\rangle$ of the ReLU network $N$. Now, we describe a polynomial space algorithm for counting the number of affine regions.
By Lemma˜A.3, the encoding size of any coefficient of an affine function that occurs in one of the affine regions of $N$ is bounded by $M:=36d^{2}n_{\max}^{2}\langle A_{\max}\rangle$, which is polynomial in $\langle N\rangle$.
As each coefficient of an affine function of $N$ is a fraction, $M$ is also an upper bound on the encoding size of a numerator and on the encoding size of a denominator. Since each affine function of $N$ is defined by $n+1$ fractions, we can exhaustively search through all sequences of $n+1$ fractions, where the numerator and denominator of each fraction can have encoding size of at most $M$. For each sequence of $n+1$ fractions, by Lemma˜B.4 we can compute in space that is polynomial in $\langle N\rangle$ if the corresponding affine function is the function of an affine region of $N$. If an affine function of an affine region is found, we increase a counter by 1. To avoid counting the same affine function more than once, we only check fraction sequences win which the numerator and denominator of every fraction are relatively prime. ∎
@@ C.1 A closed connected region which is not a closure of a union of a set of activation regions
We turn this function into a ReLU network $N$ with three hidden layers, as illustrated in Figure˜5. For the orange closed connected region $P$ in Figure˜5, there is no set of activation regions such that $P$ is the closure of a union of activation regions of the ReLU network $N$.
Figure 5: A ReLU network $N$ computing $\min(y,\max(-1,-x),\max(3-2x,-x))$. An activation region of $N$ is either a blue line, blue point, or a full dimensional cell as defined by the blue lines. There are four closed connected region as indicated by the colors. The line between the points (1,1)(1,-1) and (3,3)(3,-3) belongs to the green as well as the orange region.
@@ C.2 Further examples
Figure 6: The functions $g_{\varphi}(x)=1-\max(0,x_{1})-\max(0,1-x_{1}-x_{2})$ (top), $h_{\varphi,\varepsilon}(x)=\max(0,\varepsilon-1+g_{\varepsilon}(x))$ (center), and $f_{N_{\varphi}}(x)=\max(0,T_{2}(x)+\varepsilon-1+g_{\varphi}(x))$ (bottom) for $\varepsilon=0.4$. The function $f_{N_{\varphi}}$ is only nonzero in the blue region, which is contained in the $\varepsilon$-square (orange) around the only satisfying point of $\varphi$ (black).